# From Plausibility to Verifiability: Risk-Controlled Generative OCR for Vision-Language Models

### 20 Mar 2026 v1

@@ Abstract.
Modern vision-language models (VLMs) can act as generative OCR engines, yet open-ended decoding can expose rare but consequential failures. We identify a core deployment misalignment in generative OCR. Autoregressive decoding favors semantic plausibility, whereas OCR requires outputs that are visually grounded and geometrically verifiable. This mismatch produces severe errors, especially over-generation and unsupported substitutions, creating deployment risk even when benchmark accuracy remains high. We therefore formulate frozen VLM OCR as a selective accept/abstain problem and propose a model-agnostic Geometric Risk Controller. The controller probes multiple structured views of the same input, applies lightweight structural screening, and accepts a transcription only when cross-view consensus and stability satisfy predefined criteria, yielding a small family of operating points. Experiments on frozen VLM backbones and standard OCR benchmarks show consistent reductions in extreme-error risk and catastrophic over-generation at predictable coverage costs. Reliable deployment of generative OCR with frozen VLMs benefits from explicit system-level risk control rather than unconstrained generation.
@@ 1. Introduction
Figure 1. System overview of the proposed Geometric Risk Controller (GRC). Given multi-view queries to a frozen vision-language model, the controller applies structural screening and cross-view consensus to produce a candidate string $s^{*}$, where $s^{*}$ is the consensus transcription across views. The operating point then decides whether to accept $s^{*}$ or abstain.
@@
System overview of the Geometric Risk Controller for deployment-controlled generative OCR. An input crop is expanded into K geometric views, which are queried with a frozen vision-language model backbone. The returned transcriptions enter a controller layer with preprocessing, structural screening, cross-view consensus, and an operating-point controller. The final decision is accept or abstain, yielding either an accepted transcription or no output.
@@
Optical character recognition (OCR) is increasingly treated as an emergent capability of modern vision-language models (VLMs) (Lv et al., 2024; Yu et al., 2024). In practice, this has encouraged the use of frozen VLMs as direct generative OCR engines. However, such usage remains fragile in deployment. Under open-ended autoregressive decoding, a model may generate continuations that extend beyond the visible text or produce semantically plausible yet visually unsupported substitutions. These failure modes violate a key requirement for user-facing OCR, namely geometric verifiability, meaning that the emitted text should be locally supportable by image evidence under geometric constraints.
@@
Traditionally, OCR systems are evaluated by average-case accuracy, namely whether the predicted string matches the ground-truth transcription on a benchmark (Karatzas et al., 2013). Here the focus shifts to deployment risk, defined by the risk exposed on the outputs returned to users under a fixed inference protocol. Under this view, generative OCR with frozen VLMs becomes a selective prediction problem with abstention, where the system either emits a transcription or abstains, and abstention is part of the deployment contract rather than a convenience feature (Geifman and El-Yaniv, 2017; Wen et al., 2025). Making the accept-versus-abstain decision explicit turns evaluation from a single average-case score into an auditable risk–coverage trade-off under fixed protocols (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024). The most harmful OCR failures are both rare and highly sensitive to the inference protocol, especially the prompt template and decoding strategy. Open-ended generative decoding is optimized for coherent continuation, so when visual evidence is weak or ambiguous, the model may fall back on language priors and produce semantically plausible but visually unsupported outputs (Li et al., 2023). Consequently, stronger average benchmark performance does not necessarily reduce the severe-error tail that drives user-facing exposure. Reliability also should not depend on ad hoc prompting, since even minor prompt changes can materially alter model behavior (Zhao et al., 2021; Rudman et al., 2026). Decoding-time heuristics can likewise change hallucination frequency, making behavior harder to reproduce and audit even under a nominally fixed deployment setup (Huang et al., 2024b). Our target setting is deployment-driven. Foundation models are often shared across tasks and products, making per-application retraining for OCR costly or impractical. We therefore seek an external control layer that operates at inference time and reduces exposed risk without modifying model internals. This layer is not an after-the-fact heuristic filter; rather, the accept/abstain rule is specified as part of the deployment contract. We therefore propose a deployment-centric risk-control framework for generative OCR with frozen VLMs. At its core is the Geometric Risk Controller (GRC), which converts open-ended decoding into a selective system under a fixed accept/abstain contract. Under a fixed multi-view protocol with $K$ geometrically related views, GRC probes the frozen model through controlled input variations, applies lightweight structural validity checks, and accepts a transcription only when cross-view agreement is sufficiently stable. A single strictness knob $m$ exposes a small family of operating points on a risk–coverage curve, enabling explicit operating-point control rather than ad hoc retuning (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024). We provide full protocol details and component ablations, and report results at fixed operating points so that trade-offs remain comparable across models and datasets. While related in spirit to recent test-time scaling strategies for compact VLMs based on augmentation and aggregation (Kaya et al., 2026), our formulation is specialized to OCR failure modes and instantiated as a fixed, deployment-auditable accept/abstain contract. Our contributions are threefold:
@@
(1) Reframe. We recast frozen VLM-based OCR as a deployment-control problem, introducing geometric verifiability and deployment-oriented risk primitives that expose long-tail and catastrophic failures beyond mean accuracy.
@@
(2) Control. We propose a model-agnostic Geometric Risk Controller that converts open-ended generative OCR into a fixed-protocol selective system with an explicit strictness knob.
@@
(3) Validate. Across multiple frozen VLM backbones and scene-text benchmarks, we show that the controller consistently suppresses exposed long-tail risk while offering controllable coverage trade-offs; a compact query-budget analysis further illustrates the associated inference-cost trade-off.
@@ 2. Related Work
2.1. Generative OCR Backbones and the Deployment Misalignment
Frozen VLMs as Generative OCR Engines. Vision-language models (VLMs) are increasingly used as OCR engines (Kim et al., 2022). Systems such as KOSMOS-2.5 (Lv et al., 2024), TextHawk (Yu et al., 2024), and OCEAN-OCR (Chen et al., 2025) show strong end-to-end transcription ability across several settings. However, stronger OCR capability does not by itself guarantee deployment reliability. Under open-ended generation, autoregressive decoding optimizes next-token likelihood rather than explicit visual verifiability, leading to unsupported substitutions and runaway continuations (Huang et al., 2024b). These failures are sensitive to prompting and decoding choices (Rudman et al., 2026; Schmalfuss et al., 2025; Huang et al., 2024b), complicating auditing and control in deployment. Recent analyses further link such behaviors to broader visual-grounding weaknesses in multimodal LLMs (Tong et al., 2024).
Average-Case Metrics versus Deployment Exposure. Benchmarks such as the ICDAR series (Karatzas et al., 2013, 2015; Huang et al., 2019), OCRBench (Liu et al., 2024; Fu et al., 2025), and SVT (Wang et al., 2011) emphasize aggregated scores such as recognition accuracy, CER, and exact match. While these metrics capture average capability, they can understate rare but high-cost tail failures (Amit et al., 2025). For generative OCR, this creates a deployment misalignment between benchmark capability and exposed-output reliability, motivating an explicit accept–abstain formulation (Mao et al., 2025) under selective prediction and risk–coverage principles (Geifman and El-Yaniv, 2017).
2.2. Selective Deployment and Risk-Controlled Decision Interfaces
Selective deployment requires an explicit accept–abstain interface, not merely a stronger recognizer. Selective prediction formalizes this interface through the coverage–risk trade-off (Geifman and El-Yaniv, 2017; Wen et al., 2025; Chow, 1970), while Conformal Risk Control (CRC) formalizes finite-sample control of expected risk under a specified loss (Wang et al., 2025; Angelopoulos et al., 2024; Angelopoulos and Bates, 2021). In generative systems, however, abstention is often based on internal confidence or self-evaluation (Ren et al., 2023; Mao et al., 2025), and such signals can be prompt- and protocol-sensitive (Schmalfuss et al., 2025) while remaining overconfident on hallucinated outputs (Zollicoffer et al., 2025; Mao et al., 2025; Woo et al., 2025). Our controller adopts the same deployment-level framing but remains empirical rather than guarantee-providing, basing abstention on protocol-induced exogenous evidence—cross-view consistency and structural stability—to define auditable fixed-protocol operating points.
@@ 2.3. Verification Signals under Fixed Inference Protocols
Test-time scaling methods such as test-time augmentation, multi-crop inference, and self-consistency are widely used to improve performance and robustness through aggregation, reranking, or voting (Snell et al., 2024; Lee et al., 2025; Wang et al., 2023; Kaya et al., 2026; Chou et al., 2025a, b). Our setting uses repeated inference for a different purpose. Under a fixed protocol, cross-view agreement and stability serve as decision evidence for accept–abstain control rather than as heuristics for improving accuracy (Khan and Fu, 2024). Related OCR and structured extraction systems also use lightweight output constraints, such as character-set, length, and format checks, to suppress obviously invalid generations (Willard and Louf, 2023), but these string-level filters do not directly establish visual support (Li et al., 2025) and can remain misleading when the model is consistently wrong (Huang et al., 2024a). We therefore treat cross-view consistency under controlled geometric perturbations as exogenous evidence of visual support, turning test-time consistency into a fixed-protocol control signal for auditable accept–abstain decisions.
@@ 3. Methods
3.1. Setting and Control Contract
We study generative OCR with a frozen vision-language model (VLM) used as a black-box generator. Given an image crop $I\in\mathcal{I}$, a fixed prompt template $p$, and a fixed deterministic decoding configuration $\delta$, the frozen backbone returns
$y=G_{\delta}(I,p;\theta)\in\mathcal{Y},\qquad\theta\ \text{frozen}.$
We treat $G_{\delta}$ as the black-box mapping induced by the frozen backbone under the fixed prompt-and-decoding setup; model internals are not used, and control relies only on queried image views and returned strings.
Our goal is not to modify the backbone, but to place it inside a selective accept–abstain contract. Accordingly, the deployed OCR system is
@@ 3.1. Setting and Control Contract
where $\bot$ denotes abstention. The map $f$ is induced by an inference-time evidence protocol $\pi$ and an acceptance rule applied to that evidence.
Specifically, a protocol $\pi$ extracts from the input crop a candidate transcription $\hat{y}(I)\in\mathcal{Y}$ together with an evidence summary $e(I)$. An acceptance rule $A$ then determines whether the candidate is exposed:
$f_{\pi}(I)=\begin{cases}\hat{y}(I),&\text{if }A\!\big(e(I)\big)=1,\\ \bot,&\text{otherwise.}\end{cases}$
@@
3.2. Evidence Protocol: Multi-view Geometric Probing
We instantiate $\pi$ through multi-view probing of the same input crop. Let $\mathcal{T}$ denote a family of admissible image transforms that apply mild geometric perturbations, such as small translations, crop jitter, or scale variation, while preserving nominal text content.
A fixed protocol $\pi$ specifies the transform family $\mathcal{T}$, a view-selection rule, and the total number of queried views $K$. The original crop is always included as the anchor view,
$T_{1}=\mathrm{Id},\qquad I_{1}=I.$
The remaining views are
@@ 3.2. Evidence Protocol: Multi-view Geometric Probing
where the selection rule for $\{T_{k}\}_{k=2}^{K}$ is part of the protocol. In our implementation, the non-anchor transforms are sampled from a fixed distribution over admissible transforms. Each view is queried with the same prompt template and decoding configuration as in Sec. 3.1:
$y_{k}=G_{\delta}(I_{k},p;\theta),\qquad k=1,\dots,K.$
This yields the evidence record
$R_{\pi}(I)=\{(I_{k},y_{k})\}_{k=1}^{K}.$
Under the fixed protocol, we treat cross-view stability as an operational proxy for geometric support and cross-view disagreement as a risk signal. These exogenous signals are then used for accept–abstain control.
Figure 2. Representative accept–abstain evidence patterns under the fixed K=5K{=}5 protocol at the default operating point (m=3m{=}3). Fragmented valid outputs cause abstention, a majority can still be rejected under high dispersion, and borderline consensus is accepted when the valid views remain tightly clustered.
3.3. Canonicalization and Structural Screening
To make cross-view evidence comparable under the fixed protocol, we first map each returned string to a canonical form
@@ 3.3. Canonicalization and Structural Screening
where $\mathrm{Norm}(\cdot)$ is a protocol-fixed normalization rule that removes only presentation variance, such as whitespace normalization and, when specified by the task protocol, case canonicalization, without changing the underlying transcription content.
We then apply a structural screening rule
@@
where $\mathcal{V}\subseteq\mathcal{Y}$ is a protocol-fixed set of lightweight admissibility constraints. These constraints are label-agnostic and serve only to reject degenerate continuations, not to recognize text or estimate confidence.
The main admissibility constraint in $\mathcal{V}$ is a conservative geometric length bound. For each view $I_{k}$, we compute an upper bound $L_{\mathrm{geom}}(I_{k})$ from foreground geometry after scale normalization and dominant-axis span estimation. A decoded string is marked invalid if
@@
This bound is not intended to estimate the exact ground-truth length; it only suppresses implausibly overlong continuations.
@@
Let
$\Omega(I)=\{k:v_{k}=1\},\qquad n(I)=|\Omega(I)|.$
All downstream evidence statistics are computed only over $\Omega(I)$. This induces a minimum-evidence gate with threshold $K_{\min}$.
3.4. Decision Functional: Consensus, Stability, and Operating-Point Control
For samples with $n(I)\geq K_{\min}$, let $s^{*}(I)$ denote the unique mode of the valid canonicalized outputs $\{\tilde{y}_{k}:k\in\Omega(I)\}$ when it exists. If the mode is not unique, the system abstains. Otherwise, the candidate transcription is
$\hat{y}_{\pi}(I)=s^{*}(I),$
with vote fraction
$q(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}\mathbb{I}\!\left[\tilde{y}_{k}=s^{*}(I)\right]\in[0,1].$
To measure residual disagreement around the candidate, we use the bounded normalized edit distance

### 15 Apr 2026 v3

# From Plausibility to Verifiability: Risk-Controlled Generative OCR with Vision-Language Models

@@ Abstract.
Modern vision-language models (VLMs) can act as generative optical character recognition (OCR) engines, yet open-ended decoding can expose rare but consequential failures. We identify a core deployment misalignment in generative OCR, where autoregressive decoding favors semantic plausibility while OCR requires outputs that are visually grounded and geometrically verifiable. This mismatch produces severe errors, especially over-generation and unsupported substitutions, creating deployment risk even when benchmark accuracy remains high. We therefore formulate frozen VLM OCR as a selective accept/abstain problem and propose a model-agnostic geometric risk controller. This controller probes multiple structured views of the same input, applies lightweight structural screening, and accepts a transcription only when cross-view consensus and stability satisfy predefined criteria, yielding a small family of operating points. Experiments on frozen VLM backbones and standard OCR benchmarks show consistent reductions in extreme-error risk and catastrophic over-generation at predictable coverage costs. These results suggest that reliable deployment of generative OCR with frozen VLMs benefits from explicit system-level risk control rather than unconstrained generation. Code is available at https://github.com/phare111/GRC.
Generative OCR, Vision-Language Models, Selective Prediction, Risk Control, Abstention
@@ 1. Introduction
Figure 1. Layered system view of the proposed geometric risk controller (GRC). The pipeline consists of an input layer, a perception layer, a controller layer under protocol $\pi$, and a selective interface. A frozen vision-language model produces view-wise transcriptions in the perception layer, which are filtered and aggregated in the controller layer to determine whether the system returns a transcription or abstains.
@@
Layered architecture of the geometric risk controller for risk-controlled generative OCR. The figure contains four stages from left to right: an input layer, a perception layer, a controller layer, and a selective interface. An input crop is transformed into \$K\$ geometric views in the perception layer and queried with a frozen vision-language model backbone. The returned transcriptions are then processed in the controller layer through preprocessing, structural screening, cross-view consensus, and operating-point control. The selective interface finally either returns an accepted transcription or abstains with no output.
@@
Optical character recognition (OCR) is increasingly treated as an emergent capability of modern vision-language models (VLMs) (Lv et al., 2024; Yu et al., 2024), which has encouraged the use of frozen VLMs as direct generative OCR engines. However, open-ended autoregressive decoding remains fragile in deployment: models may over-generate beyond the visible text or produce semantically plausible but visually unsupported substitutions. Such failures violate a core requirement of user-facing OCR, namely geometric verifiability, meaning that emitted text should remain locally supportable by image evidence under geometric constraints.
Rather than focusing only on average-case benchmark accuracy (Karatzas et al., 2013), we study deployment risk, that is, the risk exposed on outputs returned to users under a fixed inference protocol. Under this view, frozen-VLM OCR becomes a selective prediction problem with abstention: the system either emits a transcription or abstains, and abstention is part of the deployment contract rather than a convenience feature (Geifman and El-Yaniv, 2017; Wen et al., 2025). This shifts evaluation from a single accuracy number to an auditable risk–coverage trade-off under fixed protocols (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024).
The severe-error tail is especially sensitive to prompting and decoding. When visual evidence is weak or ambiguous, open-ended generation may fall back on language priors and produce visually unsupported but semantically plausible outputs (Li et al., 2023). As a result, stronger benchmark accuracy does not necessarily reduce user-facing catastrophic failures. Reliability should also not depend on ad hoc prompt or decoding choices, since even small changes can materially alter behavior and undermine reproducibility (Zhao et al., 2021; Rudman et al., 2026; Huang et al., 2024b).
@@
We therefore propose a deployment-centric risk-control framework for generative OCR with frozen VLMs. Its core component, the geometric risk controller (GRC), converts open-ended decoding into a selective system under a fixed accept/abstain contract. Under a fixed multi-view protocol with $K$ geometrically related views, GRC probes the frozen model through controlled input variations, applies lightweight structural validity checks, and accepts a transcription only when cross-view agreement is sufficiently stable. A single strictness knob $m$ exposes a small family of operating points on a risk–coverage curve, enabling explicit operating-point control rather than ad hoc retuning (Geifman and El-Yaniv, 2017; Angelopoulos et al., 2024). We report fixed-protocol results, full component ablations, and compact query-budget analyses so that trade-offs remain comparable across models and datasets. While related in spirit to recent test-time scaling strategies based on augmentation and aggregation (Kaya et al., 2026), our formulation is specialized to OCR failure modes and instantiated as a fixed, deployment-auditable accept/abstain contract.
Our contributions are threefold:
@@
(1) We recast frozen VLM-based OCR as a deployment-control problem, introducing geometric verifiability and deployment-oriented risk primitives that reveal long-tail and catastrophic failures beyond mean accuracy.
@@
(2) We develop a model-agnostic geometric risk controller (GRC), converting open-ended generative OCR into a fixed-protocol selective system with explicit operating points and a controllable strictness knob.
@@
(3) We demonstrate across multiple frozen VLM backbones and scene-text benchmarks that GRC consistently suppresses exposed long-tail risk while providing controllable coverage trade-offs and a compact view of the associated inference-cost trade-off.
@@ 2. Related Work
2.1. Generative OCR Backbones and Deployment Misalignment
Frozen VLMs as Generative OCR Engines. VLMs are increasingly used as OCR engines (Kim et al., 2022). Systems such as KOSMOS-2.5 (Lv et al., 2024), TextHawk (Yu et al., 2024), and OCEAN-OCR (Chen et al., 2025) show strong end-to-end transcription ability across several settings. However, stronger OCR capability does not by itself guarantee deployment reliability. Under open-ended generation, autoregressive decoding optimizes next-token likelihood rather than explicit visual verifiability, leading to unsupported substitutions and runaway continuations (Huang et al., 2024b). These failures are sensitive to prompting and decoding choices (Rudman et al., 2026; Schmalfuss et al., 2025; Huang et al., 2024b), complicating auditing and control in deployment. Recent analyses further link such behaviors to broader visual-grounding weaknesses in multimodal LLMs (Tong et al., 2024).
Average-Case Metrics versus Deployment Exposure. Benchmarks such as the ICDAR series (Karatzas et al., 2013, 2015; Huang et al., 2019), OCRBench (Liu et al., 2024; Fu et al., 2025), and SVT (Wang et al., 2011) emphasize aggregated scores such as recognition accuracy, CER, and exact match. While these metrics capture average capability, they can understate rare but high-cost tail failures (Amit et al., 2025). For generative OCR, this creates a deployment misalignment between benchmark capability and exposed-output reliability, motivating an explicit accept/abstain formulation (Mao et al., 2025) under selective prediction and risk–coverage principles (Geifman and El-Yaniv, 2017).
$2.$ Selective Deployment and Risk-Controlled Decision Interfaces
Selective deployment requires an explicit accept/abstain interface, not merely a stronger recognizer. Selective prediction formalizes this interface through the coverage–risk trade-off (Geifman and El-Yaniv, 2017; Wen et al., 2025; Chow, 1970), while Conformal Risk Control (CRC) formalizes finite-sample control of expected risk under a specified loss (Wang et al., 2025; Angelopoulos et al., 2024; Angelopoulos and Bates, 2021). In generative systems, however, abstention is often based on internal confidence or self-evaluation (Ren et al., 2023; Mao et al., 2025), and such signals can be prompt- and protocol-sensitive (Schmalfuss et al., 2025) while remaining overconfident on hallucinated outputs (Zollicoffer et al., 2025; Mao et al., 2025; Woo et al., 2025). Our controller adopts the same deployment-level framing but remains empirical rather than guarantee-providing, basing abstention on protocol-induced exogenous evidence—cross-view consistency and structural stability—to define auditable fixed-protocol operating points.
@@ 2.3. Verification Signals under Fixed Inference Protocols
Test-time scaling methods such as test-time augmentation, multi-crop inference, and self-consistency are widely used to improve performance and robustness through aggregation, reranking, or voting (Snell et al., 2024; Lee et al., 2025; Wang et al., 2023; Kaya et al., 2026; Chou et al., 2025a, b). Our setting uses repeated inference for a different purpose. Under a fixed protocol, cross-view agreement and stability serve as decision evidence for accept/abstain control rather than as heuristics for improving accuracy (Khan and Fu, 2024). Related OCR and structured extraction systems also use lightweight output constraints, such as character-set, length, and format checks, to suppress obviously invalid generations (Willard and Louf, 2023), but these string-level filters do not directly establish visual support (Li et al., 2025) and can remain misleading when the model is consistently wrong (Huang et al., 2024a). We therefore treat cross-view consistency under controlled geometric perturbations as exogenous evidence of visual support, turning test-time consistency into a fixed-protocol control signal for auditable accept/abstain decisions.
@@ 3. Methods
Figure 1 provides an overview of the proposed geometric risk controller (GRC). Given an input crop II, the fixed protocol $\pi$ generates multiple geometric views $\{I_{k}\}_{k=1}^{K}$, which are queried with the frozen generator to produce raw outputs $\{y_{k}\}_{k=1}^{K}$. The controller then canonicalizes these outputs, applies structural screening to obtain the valid set $\Omega(I)$, extracts a consensus candidate $s^{*}(I)$ together with summary statistics such as the vote fraction $q(I)$ and dispersion $\Delta(I)$, and finally applies the operating-point rule to either expose the candidate or abstain.
3.1. Problem Setting and Control Contract
We study generative OCR with a frozen VLM used as a black-box generator. Given an input crop $I\in\mathcal{I}$, a fixed prompt template $p$, and a fixed deterministic decoding configuration $\delta$, the backbone with frozen parameters $\theta$ returns a decoded string
$y=G_{\theta,\delta}(I,p)\in\mathcal{Y},$
where $\mathcal{I}$ is the space of input crops and $\mathcal{Y}$ is the space of decoded strings. Both $\theta$ and $\delta$ are fixed throughout deployment. Thus, $G_{\theta,\delta}$ is treated as a black-box mapping under the fixed prompt-and-decoding setup. Control bypasses model internals, relying only on queried image views and returned strings.
Our goal is not to modify the backbone, but to place it inside a selective accept/abstain contract. Accordingly, the deployed OCR system is a map
@@ 3.1. Problem Setting and Control Contract
where $\bot$ denotes abstention. The system output is induced by an inference-time evidence protocol $\pi$ together with an acceptance rule applied to the resulting evidence.
Specifically, for an input crop II, protocol $\pi$ produces a candidate transcription $\hat{y}_{\pi}(I)\in\mathcal{Y}$ and an evidence summary $e_{\pi}(I)$. An acceptance rule AA then determines whether the candidate is exposed:
$f_{\pi}(I)=\begin{cases}\hat{y}_{\pi}(I),&\text{if }A\!\left(e_{\pi}(I)\right)=1,\\ \bot,&\text{otherwise.}\end{cases}$
@@
3.2. Evidence Protocol and Multi-View Geometric Probing
We instantiate protocol $\pi$ by querying multiple geometric views of the same input crop II. Let $\mathcal{T}$ denote a family of admissible image transforms, where each $T\in\mathcal{T}$ applies a mild geometric perturbation—such as a small translation, crop jitter, or scale variation—while preserving the nominal text content.
A fixed protocol $\pi$ specifies the transform family $\mathcal{T}$, a rule for selecting concrete transforms, and the total number of queried views $K$. The original crop is always included as the anchor view,
$T_{1}=\mathrm{\textit{Id}},\qquad I_{1}=I,$
where $\mathrm{Id}$ is the identity transform. The remaining views are
@@ 3.2. Evidence Protocol and Multi-View Geometric Probing
where each $T_{k}\in\mathcal{T}$ is chosen according to the protocol. Thus, II denotes the original input crop, while $I_{k}$ denotes its $k$-th transformed view.
Each view $I_{k}$ is queried with the same prompt template $p$ using the frozen generator $G_{\theta,\delta}$:
$y_{k}=G_{\theta,\delta}(I_{k},p),\qquad k=1,\dots,K.$
Here, $y_{k}$ is the raw decoded string returned for view $I_{k}$. This yields the evidence record
$R_{\pi}(I)=\{(I_{k},y_{k})\}_{k=1}^{K},$
namely, the collection of queried views and their returned strings under protocol $\pi$. Under the fixed protocol, cross-view stability serves as an operational proxy for geometric support, while cross-view disagreement acts as a risk signal for the final accept/abstain decision.
$3.$ Canonicalization and Structural Screening
To make cross-view evidence comparable, we map each raw output $y_{k}$ to a canonical form
@@
where $\mathrm{Norm}(\cdot)$ is a protocol-fixed normalization rule. It removes only presentation variance, such as whitespace normalization and, when required by the task protocol, case canonicalization, without changing the underlying transcription content.
We then apply structural screening through
@@
where $v_{k}$ is a binary validity indicator and $\mathcal{V}\subseteq\mathcal{Y}$ is a protocol-fixed set of lightweight admissibility constraints. These constraints are label-agnostic and are used only to reject degenerate continuations.
The main constraint in $\mathcal{V}$ is a conservative geometric length bound. For each view $I_{k}$, we compute an upper bound $L_{\mathrm{geom}}(I_{k})$ from foreground geometry after scale normalization and dominant-axis span estimation. Writing $|\tilde{y}_{k}|$ for the character length of $\tilde{y}_{k}$, we mark an output as invalid whenever
@@
This bound is not intended to estimate the exact ground-truth length. Rather, it serves only to suppress implausibly overlong continuations.
@@
We then define the valid-view index set
$\Omega(I)=\{k\in\{1,\dots,K\}:v_{k}=1\},$
and its cardinality
$n(I)=|\Omega(I)|.$
All downstream evidence statistics are computed only over the valid set $\Omega(I)$. This induces a minimum-evidence gate with threshold $K_{\min}$.
3.4. Decision Functional and Operating-Point Control
Figure 1 sketches the key quantities formalized below: the consensus candidate $s^{*}(I)$, its support $q(I)$, the dispersion $\Delta(I)$, and the operating-point thresholding rule. For inputs with $n(I)\geq K_{\min}$, let $s^{*}(I)$ denote the unique mode of the valid canonicalized outputs $\{\tilde{y}_{k}:k\in\Omega(I)\}$, when such a mode exists. If no unique mode exists, the controller abstains. Otherwise, the candidate transcription is
$\hat{y}_{\pi}(I)=s^{*}(I).$
To quantify support for the candidate, we define the vote fraction
$q(I)=\frac{1}{n(I)}\sum_{k\in\Omega(I)}\mathbb{I}\!\left[\tilde{y}_{k}=s^{*}(I)\right]\in[0,1],$
that is, the fraction of valid views whose canonicalized outputs exactly match $s^{*}(I)$.
To quantify residual disagreement around the candidate, we define the bounded normalized edit distance